OpenStudy (jacky90):

The time it takes to cover the distance between two cities by car varies inversely with the speed of the car. The trip takes 4 hours for a car moving at 50 mph . How long does the trip take for a car moving at 40 mph?

OpenStudy (dayakar):

\[time \alpha \frac{ 1 }{speed}\] s2*t2 = s1*t1 s2 =40mph, t2= ?, s1= 50mph,t1=4hr can u complete remaining part

OpenStudy (jacky90):

Im stumped how would input this?

OpenStudy (jacky90):

Would I*

OpenStudy (dayakar):

s2*t2 = s1 *t1 40 *t2 = 50 *4 \[t2 =\frac{ 50 \times 4 }{ 40 }\] find t2 now

OpenStudy (dayakar):

is it clear

OpenStudy (dayakar):

plz, respond, do u understand or not

OpenStudy (jacky90):

I understood, thanks for that last answer.